# Tuning Algorithmic and Architectural Hyperparameters in Graph-Based Semi-Supervised Learning with Provable Guarantees

### v1

@@ Appendix C Proofs in Section 4
$\left\|h_{i}^{L}(\eta,U,V)-h_{i}^{L}(\eta^{\prime},U^{\prime},V^{\prime})\right\|$
$\leq$ $C_{U}\left(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1}\right)|\eta-\eta^{\prime}|+rC_{U}\left(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1}\right)+\left(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1}\right)\|U-U^{\prime}\|$
$+C_{U}\left(\max_{j\in\mathcal{N}_{i}}\left\|h_{j}^{L-1}-h_{j}^{\prime(L-1)}\right\|\right)+\frac{2rC_{U}}{C_{dl}}\left\|h_{i}^{L-1}-h_{i}^{\prime(L-1)}\right\|+\frac{2r}{C_{dl}}\bar{h}_{i}^{L-1}\|U-U^{\prime}\|+\frac{2rC_{U}}{C_{dl}}|\eta-\eta^{\prime}|$
$\leq$ $k_{1}+k_{2}|\eta-\eta^{\prime}|+k_{3}\|U-U^{\prime}\|+k_{4}\left(\max_{j\in[n]}\|h_{j}^{L-1}-h_{j}^{\prime(L-1)}\|\right)$
$=$ $\left(k_{1}+k_{2}|\eta-\eta^{\prime}|+k_{3}\|U-U^{\prime}\|\right)\frac{k_{4}^{L}-k_{4}}{k_{4}-1}+k_{4}\left(\max_{j\in[n]}\|h^{0}_{j}-h_{j}^{\prime 0}\|\right)$
$\leq$ $\left(k_{1}+k_{2}|\eta-\eta^{\prime}|+k_{3}\|U-U^{\prime}\|\right)\frac{k_{4}^{L}-k_{4}}{k_{4}-1}+k_{4}C_{z}\|U-U^{\prime}\|$
@@
$k_{1}$ $=r^{L}C_{U}^{L+1}C_{z}\max\left(1,\frac{1}{C_{dl}}\right)^{L-1}$
$k_{3}$ $=\left(1+\frac{2r}{C_{dl}}\right)r^{L-1}C_{U}^{L}C_{z}\max\left(1,\frac{1}{C_{dl}}\right)^{L-1}$
@@
$\mathcal{N}\left(\eta,\frac{\epsilon}{2A},|\cdot|\right)\leq 1+\frac{4A}{\epsilon}.$
@@
$\mathcal{N}\left(U,\frac{\epsilon}{2B},\|\cdot\|_{F}\right)\leq\left(1+\frac{4BC_{U}\sqrt{d}}{\epsilon}\right)^{d^{2}}.$
@@
$P\leq$ $\mathcal{N}\left(\eta,\frac{\epsilon}{2A},|\cdot|\right)\cdot\mathcal{N}\left(U,\frac{\epsilon}{2B},\|\cdot\|_{F}\right)$
$\leq$ $\left(1+\frac{4A}{\epsilon}\right)\left(1+\frac{4BC_{U}\sqrt{d}}{\epsilon}\right)^{d^{2}}\leq\left(1+\frac{4\max\{A,BC_{U}\sqrt{d}\}}{\epsilon}\right)^{d^{2}+1}$
@@
$\hat{R}_{\mathcal{T}}\left(\mathcal{H}^{\gamma}_{(\eta,U,V))}\right)\leq\frac{4}{m}+\frac{12\sqrt{(d^{2}+1)\log\left(8\sqrt{m}\max\{A,BC_{U}\sqrt{d}\}\right)}}{\sqrt{m}}.$
@@ Appendix D Experiments
We present here additional details related to our empirical studies.
@@ D.3 Experiment Setup for GCAN
These GCAN interpolation experiments are all run with only $20\%$ of the dataset being labeled datapoints, and the remaining $80\%$ representing the unlabeled datapoints that we test our classification accuracy on. Table 3 notes the exact setup of each dataset, and the overall training time of each experiment.

### v2

@@ Appendix C Proofs in Section 4
$\|h_{i}^{L}(\eta,U,V)-h_{i}^{L}(\eta^{\prime},U^{\prime},V^{\prime})\|$
$\qquad\leq C_{U}(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1})|\eta-\eta^{\prime}|+rC_{U}(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1})+(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1})\|U-U^{\prime}\|$
$\qquad\qquad+C_{U}(\max_{j\in\mathcal{N}_{i}}\|h_{j}^{L-1}-h_{j}^{\prime(L-1)}\|)+\frac{2rC_{U}}{C_{dl}}\|h_{i}^{L-1}-h_{i}^{\prime(L-1)}\|+\frac{2r}{C_{dl}}\bar{h}_{i}^{L-1}\|U-U^{\prime}\|+\frac{2rC_{U}}{C_{dl}}|\eta-\eta^{\prime}|$
$\qquad\leq k_{1}+k_{2}|\eta-\eta^{\prime}|+k_{3}\|U-U^{\prime}\|+k_{4}(\max_{j\in[n]}\|h_{j}^{L-1}-h_{j}^{\prime(L-1)}\|)$
$\qquad=\left(k_{1}+k_{2}|\eta-\eta^{\prime}|+k_{3}\|U-U^{\prime}\|\right)\frac{k_{4}^{L}-k_{4}}{k_{4}-1}+k_{4}(\max_{j\in[n]}\|h^{0}_{j}-h_{j}^{\prime 0}\|)$
$\qquad\leq\left(k_{1}+k_{2}|\eta-\eta^{\prime}|+k_{3}\|U-U^{\prime}\|\right)\frac{k_{4}^{L}-k_{4}}{k_{4}-1}+k_{4}C_{z}\|U-U^{\prime}\|$
@@
$k_{1}$ $=r^{L}C_{U}^{L+1}C_{z}\max(1,\frac{1}{C_{dl}})^{L-1}$
$k_{2}$ $=r^{L-1}C_{U}^{L+1}C_{z}\max(1,\frac{1}{C_{dl}})^{L-1}+\frac{2rC_{U}}{C_{dl}}$
$k_{3}$ $=\left(1+\frac{2r}{C_{dl}}\right)r^{L-1}C_{U}^{L}C_{z}\max(1,\frac{1}{C_{dl}})^{L-1}$
@@
$\mathcal{N}(\eta,\frac{\epsilon}{2A},|\cdot|)\leq 1+\frac{4A}{\epsilon}.$
@@
$\mathcal{N}(U,\frac{\epsilon}{2B},\|\cdot\|_{F})\leq\left(1+\frac{4BC_{U}\sqrt{d}}{\epsilon}\right)^{d^{2}}.$
@@
$P\leq$ $\mathcal{N}(\eta,\frac{\epsilon}{2A},|\cdot|)\cdot\mathcal{N}(U,\frac{\epsilon}{2B},\|\cdot\|_{F})$
$\leq$ $\left(1+\frac{4A}{\epsilon}\right)\left(1+\frac{4BC_{U}\sqrt{d}}{\epsilon}\right)^{d^{2}}\leq(1+\frac{4\max\{A,BC_{U}\sqrt{d}\}}{\epsilon})^{d^{2}+1}$
@@
$\hat{R}_{\mathcal{T}}(\mathcal{H}^{\gamma}_{(\eta,U,V))})\leq\frac{4}{m}+\frac{12\sqrt{(d^{2}+1)\log\left(8\sqrt{m}\max\{A,BC_{U}\sqrt{d}\}\right)}}{\sqrt{m}}.$
@@ D.3 Experiment Setup for GCAN
These GCAN interpolation experiments are all run with only $20\%$ of the dataset being labeled datapoints, and the remaining $80\%$ representing the unlabeled datapoints that we test our classification accuracy on. Table 3 notes the exact setup of each dataset, and the overall training time of each experiment. We would like to examine our theory with the simplest network that is still non-linear, so we selected a hidden dimension being 1. Note that our theory on sample complexity bounds still applies to larger networks, but implementing our techniques on larger networks and larger graphs might require additional computational improvements.